I need help with this problem. I do not know where to begin. Can someone please help me. Thank you. This is from a worksheet.

Write a formula for the nth term of the infinite sequence whose nth term is given.

7.) -4,-2,0,2,4...

Answer by stanbon(75887)

-4,-2,0,2,4...
----------------
a(1) = -4
d = -2--4 = 2
----------------
a(n) = a(1) + (n-1)d
a(n) = -4 + (n-1)*2
a(n) = -4 +2n-2
a(n) = -6 + 2n
===================
Cheers,
Stan H.